### v1

# Unavoidable Canonical Nonlinearity Induced by Gaussian Measures Discretization

Abstract
When we consider canonical averages for classical discrete systems, typically referred to as substitutional alloys, the map $\phi$ from many-body interatomic interactions to thermodynamic equilibrium configurations generally exhibits complicated nonlinearity. This canonical nonlinearity is fundamentally rooted in deviations of the discrete configurational density of states (CDOS) from continuous Gaussian families, and has conventionally been characterized by the Kullback-Leibler (KL) divergence on discrete statistical manifold. Thus, the previous works inevitablly missed intrinsic nonlinearities induced by discretization of Gaussian families, which remains invisible within conventional information-geometric descriptions. In the present work, we identify and quantify such unavoidable canonical nonlinearity by employing the 2-Wasserstein distance with a cost function aligned with the Fisher metric for Gaussian families. We derive an explicit expression for the Wasserstein distance in the limit of vanishing discretization scale $d\to 0$: $W_{2}=d\sqrt{\textrm{Tr}\left(\Gamma^{-1}\right)/12}$, where $\Gamma$ denotes covariance matrix of the CDOS. We further show that this limiting Wasserstein distance admits a clear geometric interpretation on the statistical manifold, equivalent to a KL divergence associated with the expected parallel translations of continuous Gaussian. Our framework thus provides a transport-information-geometric characterization of discretization-induced nonlinearity in classical discrete systems, with future potential applications to configurational thermodynamics.
I Introduction
Canonical averages play a central role in statistical thermodynamics, providing a fundamental link between microscopic interactions and thermodynamic equilibrium configurations. For classical discrete systems with $f$ structural degrees of freedom (SDFs) on a given lattice, such as substitutional alloys, this correspondence is expressed as
$\Braket{q_{p}}_{Z}=Z^{-1}\sum_{i}q_{p}^{(i)}\exp\left(-\beta U^{(i)}\right),$ (1)
where $\{q_{1},\cdots,q_{f}\}$ denotes a complete set of structural coordinates, $\Braket{\cdot}_{Z}$ the canonical average, $\beta$ the inverse temperature, and $Z=\sum_{i}\exp\left(-\beta U^{(i)}\right)$ the partition function, with the summation taken over all microscopic configurations $i$.
When complete orthonormal basis functions (such as the generalized Ising modelce ) are adopted for the coordinates, the potential energy of configuration $k$ is given by
$U^{(k)}=\sum_{j=1}^{f}C_{j}q_{j}^{(k)},$ (2)
where the expansion coefficients are given by inner products $C_{j}=\Braket{U|q_{j}}$, i.e., trace over possible configurations.
Introducing the vectors $\mathbf{Q}_{Z}=(\Braket{q_{1}}_{Z},\cdots,\Braket{q_{f}}_{Z})$ and $\mathbf{U}=(C_{1},\cdots,C_{f})$, the canonical average in Eq. (1) defines a map
$\phi:\mathbf{U}\mapsto\mathbf{Q}_{Z}.$ (3)
It is well known that this map is generally nonlinear. Only in exceptional cases where the configurational density of states (CDOS) forms a multivariate Gaussian distribution, $\phi$ reduces to a globally linear map ig . In realistic discrete systems, however, the CDOS inevitably deviates from Gaussian families due to the discrete nature of configuration space imposed by lattice constraints.
Existing approaches to quantify such canonical nonlinearity are primarily based on information geometry, in particular the Kullback-Leibler (KL) divergence: The nonlinearity was evaluated by comparing the discrete CDOS of a real system with a reference Gaussian distribution having the same mean and covariance matrix, where the Gaussian is discretized on the same configurational support as the real CDOS. While this framework captures deviations of the CDOS from Gaussian, it necessarily includes additional contributions arising from the discretization of continuous Gaussian families themselves. Such discretization-induced effects remain invisible on discrete statistical manifolds, where probability measures are compared only on a fixed set of discrete supports.
In the present work, we identify this contribution as an unavoidable canonical nonlinearity (UCN). To quantify the UCN, a framework is required that is sensitive to probability weights as well as to the geometric rearrangement of probability mass induced by discretization. Motivated by this requirement, we employ optimal transport theory, specifically the 2-Wasserstein distance $W_{2}$. By introducing a quadratic cost function aligned with the Fisher metric of Gaussian families, the Wasserstein distance becomes directly compatible with information-geometric descriptions while remaining sensitive to discretization-induced geometric distortions. We derive an explicit expression for $W_{2}$ between a continuous Gaussian and its discretized counterpart in the limit of vanishing discretization scale $d\to 0$, yielding a universal form depending only on the covariance matrix of the CDOS and $d$. We further show that this limiting Wasserstein distance admits a clear geometric interpretation on the statistical manifold, equivalent to a KL divergence associated with the expected parallel translations of continuous Gaussian.
The present study thus provides a transport-information-geometric characterization of discretization-induced nonlinearity in classical discrete systems, which has remained inaccessible to information-geometric approaches alone. The details are shown below.
II Concept and Derivation
Unavoidable Canonical Nonliearity
Here, we briefly clarify the concept of canonical nonlinearity and its unavoidable contribution, referred to as the unavoidable canonical nonlinearity (UCN). Let $P\left(q\right)$ denote the CDOS of a realistic discrete system, given as a discrete probability distribution. We also introduce a reference continuous Gaussian distribution $P_{c}\left(q\right)$, which shares the same mean vector and covariance matrix $\Gamma$ as the realistic CDOS. It is well known that when the CDOS is exactly given by the continuous Gaussian $P_{c}$, the canonical average map $\phi$ in Eq. (3) becomesem2
$^{\forall}{C_{j}},\ \mathbf{Q}_{Z}=\left(-\beta\Gamma\right)^{-1}\cdot\mathbf{U},$ (4)
which implies that $\phi$ reduces to a globally linear map. From this viewpoint, canonical nonlinearity originates from deviations of the realistic CDOS from Gaussian families.
Based on this observation, previous studies have quantified canonical nonlinearity by measuring the difference between $P$ and $P_{c}$ using the KL divergence. Since the KL divergence requires the two distributions to be defined on the same support (or to satisfy an inclusion relationship of supports), the continuous Gaussian $P_{c}$ has been discretized on the same configurational support as $P$, yielding a discretized Gaussian distribution $P_{d}$. Accordingly, canonical nonlinearity has been characterized by the KL divergence $D\left(P:P_{d}\right)$, which extends to the KL divergence between the canonical distributions induced by the CDOSs $P$ and $P_{d}$.
While this framework successfully captures non-Gaussian features inherent to the realistic CDOS, it inevitably includes an additional contribution originating from the discretization of the continuous Gaussian distribution itself. This contribution cannot be attributed to intrinsic non-Gaussian feature of the CDOS. Instead, it reflects geometric distortion of probability mass induced by the discretization process.
We identify this contribution as the unavoidable canonical nonlinearity, UCN. By construction, the UCN arises solely from discretizing continuous Gaussian families onto a discrete configurational support. Consequently, this contribution is fundamentally inaccessible to conventional information-geometric approaches formulated on discrete statistical manifolds, which compare probability weights only on fixed discrete supports. To isolate and quantify the UCN, it is therefore necessary to compare a continuous Gaussian distribution with its discretized counterpart in a framework that satisfies the following requirements: (i) it captures not only probability weights but also the geometric rearrangement of probability mass induced by discretization, and (ii) it remains consistent with, or admits a reinterpretation within, the existing information-geometric description on statistical manifolds. These considerations naturally motivate the present use of optimal transport theory.
Wesserstein Distance Aligned with Fisher Metric
Optimal transport theory can provide a natural framework to compare probability distributions defined on different supports, e.g., continuous and discrete ones. Among various transport distances, we employ the 2-Wasserstein distance $W_{2}$, which is expected to be particularly suitable for the present purpose, since it can provide compatible description for Gaussian families on statistical manifold, through its appropriate definition of the cost function.
Let $Q$ and $R$ be probability measures on $\mathbb{R}^{f}$. Then the squared 2-Wasserstein distance is given byw2
$W_{2}^{2}\left(Q,R\right)=\inf_{\pi\in\Pi\left(Q,R\right)}\int c\left(x,y\right)\,d\pi\left(x,y\right),$ (5)
where $\Pi(P,Q)$ denotes the set of all couplings (or transport plan) with marginals $P$ and $Q$, and $c\left(x,y\right)$ denotes cost function, i.e., transport cost from $Q\left(x\right)$ to $R\left(y\right)$. To ensure the compatibility with statistical manifold, we introduce a cost function as the quadratic form weighted by the inverse covariance matrix $\Gamma^{-1}$ of Gaussian family, namely,
$c\left(x,y\right)=\left(x-y\right)^{\mathrm{T}}\Gamma^{-1}\left(x-y\right).$ (6)
This choice can be viewed as replacing the standard Euclidean squared distance $\|x-y\|^{2}$ in the definition of the conventional 2-Wasserstein distance by the squared distance induced by the Fisher metric on the Gaussian statistical manifold. When we consider Gaussian families, its landscape is completely determined by mean vector $\mu$ and covariance matrix $\Gamma$. Considering that (i) $\mu\in\mathbb{R}^{f}$ and $\Gamma\in\mathbf{M}_{f,f}\left(\mathbb{R}\right)$ and (ii) transport object is in $\mathbb{R}^{f}$, appropriate Fisher metric is associated with translations of the mean while fixing covariance matrix, which reduces to the constant metric tensor $\Gamma^{-1}$. In fact, under the definition of the cost function of Eq. (6), we see the following relationship for Gaussian families:
$^{\forall}\delta\mu\in\mathbb{R}^{f},\$$W^{2}_{2}\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right)$ (7)
$=2D\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right),$
under the restriction of fixed covariance matrix and only allowing translations of its mean.
Under these preparations, we should quantify $W_{2}\left(P_{c},P_{d}\right)$, i.e., 2-Wesserstein distance between a continuous multivariate Gaussian distribution $P_{c}$ with covariance matrix $\Gamma$, and its discretized counterpart $P_{d}$. Here, we consider discretization process as partitioning the configuration space ($\simeq\mathbb{R}^{f}$) into hypercubic cells $V_{k}$ of side length $d$, then taking the vanishing discretization limit of $d\to 0$. Explicitly, this process is expressed as
$^{\forall}k,\ P_{d}\left(q_{k}\right)=\int_{V_{k}}P_{c}\left(q\right)dq,\ d\to 0.$ (8)
In the limit $d\to 0$, the optimal transport plan is expected to converge to a local projection that maps each point $q\in V_{k}$ to the representative point $q^{\prime}_{k}$, namely,
$W_{2}^{2}\left(P_{c},P_{d}\right)=\sum_{k}\int_{V_{k}}\left(q-q^{\prime}_{k}\right)^{\textrm{T}}\Gamma^{-1}\left(q-q^{\prime}_{k}\right)P_{c}\left(q\right)dq.$ (9)
When we introduce the following variable transform
$u=q-q^{\prime}_{k},\ ^{\forall}i:\ \left|u_{i}\right|\leq\frac{d}{2}$ (10)
with representative hypercubic as $V_{0}$, and taking Tailor series expantion of $P_{c}\left(q^{\prime}_{k}+u\right)$ around $q^{\prime}_{k}$, we can reasonably retain the leading order in $d$, thereby
$W_{2}^{2}\left(P_{c}:P_{d}\right)$ $=$ $\sum_{k}P_{c}\left(q_{k}\right)\int_{V_{0}}u^{\textrm{T}}\Gamma^{-1}udu$ (11)
$=$ $\left\{\sum_{k}P_{c}\left(q_{k}\right)\right\}\left\{\sum_{i}\left(\Gamma^{-1}\right)_{ii}\frac{d^{f+2}}{12}\right\}$
$=$ $d^{2}\frac{1}{12}\textrm{Tr}\left(\Gamma^{-1}\right).$
The last equation can be obtained since at $d\to 0$,
$\sum_{k}P_{c}\left(q_{k}\right)d^{f}=\int_{\mathbb{R}^{f}}P_{c}\left(q\right)dq=1.$ (12)
Remarkably, this expression is universal: it depends only on the covariance matrix of the Gaussian distribution and the discretization scale $d$, and is independent of higher-order moments induced by the discretization process. This result provides a quantitative measure of the UCN induced purely by discretization.
Information-Geometric Intepretation of $W_{2}$
We finally provide an information-geometric interpretation of the derived Wasserstein distance in Eq. (11), clarifying its equivalence to the KL divergence associated with expected parallel translations of Gaussian distributions. Under fixed $\Gamma$, the KL between the following two Gaussian distribution is exactly given bykl
$^{\forall}\delta\mu\in\mathbb{R}^{f},\ D\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right)=\frac{1}{2}\delta\mu\Gamma^{-1}\delta\mu.$
This relation certainly shows that under fixed $\Gamma$, the KL divergence provides a well-known quadratic form with the (block-diagonal part of) Fisher metric tensor $\Gamma^{-1}$ as discussed above. To derive the $W_{2}$ distance in Eq. (11), we perform discretization of Gaussian for $f$-dimensional hypercubic with side length $d$. At the side of statistical manifold, we first extend the translation magnitude of Gaussian $\delta\mu$ to random vector as an i.i.d. random vector with probability density $\rho$ given by the uniform distribution, namely,
$\rho\left(\delta\mu\right)=\textrm{Unif}\left[-\frac{d}{2},\frac{d}{2}\right]^{f}\ \left(\textrm{i.i.d.}\right).$ (14)
Under this extended definition, we consider the following expectation:
$\mathbb{E}_{\rho}\left[\delta\mu^{\textrm{T}}\Gamma^{-1}\delta\mu\right]=\sum_{i,k}\left(\Gamma^{-1}\right)_{ik}\mathbb{E}_{\rho}\left[\delta\mu_{i}\delta\mu_{k}\right].$ (15)
Since $\mathbb{E}_{\rho}\left[\delta\mu_{i}\delta\mu_{k}\right]$ corresponds to $\left(i,k\right)$-component of the covariance matrix for $\rho\left(\delta\mu\right)$, from Eq. (14),
$\mathbb{E}_{\rho}\left[\delta\mu_{i}\delta\mu_{k}\right]_{i\neq k}=0$
$\mathbb{E}_{\rho}\left[\delta\mu_{i}\delta\mu_{i}\right]=\int_{-d/2}^{d/2}\frac{1}{d}x^{2}dx=\frac{1}{12}d^{2},$ (16)
thereby
$\mathbb{E}_{\rho}\left[\delta\mu^{\textrm{T}}\Gamma^{-1}\delta\mu\right]=d^{2}\frac{1}{12}\textrm{Tr}\left(\Gamma^{-1}\right).$ (17)
We therefore obtain the important relationships:
$\lim_{d\to 0}W_{2}^{2}\left(P_{c}:P_{d}\right)$ $=$ $d^{2}\frac{1}{12}\textrm{Tr}\left(\Gamma^{-1}\right).$
$=$ $2\cdot E_{\rho}\left[D\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right)\right].$
Therefore, the squared Wasserstein distance between the continuous Gaussian distribution and its discretized counterpart is equivalent to twice the KL divergence averaged over continuous Gaussian translations generated by the discretization procedure.
This result establishes a clear geometric interpretation of the UCN: The cost to project continuous Gaussian onto a discrete configurational support is given by the cumulative information-geometric distance for expected parallel translations within the Gaussian statistical manifold. These contributions are invisible to KL-based comparisons restricted to discrete statistical manifolds, but are naturally captured by the present transport-information-geometric framework.
III Conclusions
In this work, we investigated the intrinsic canonical nonlinearity arises from discrepancy between continuous Gaussian families and their discretized counterparts, based on optimal transport and information geometry. By focusing on the vanishing discretization limit for 2-Wasserstein distance with its cost function aligned with Fisher metric for Gaussian families, we demonstrated that the residual $W_{2}$ distance is determined solely by the inverse covariance matrix and the discretization scale. Our central result exhibits that the derived $W_{2}$ distance admits a clear information-geometric interpretation: it is exactly equivalent to twice the Kullback-Leibler divergence averaged over random parallel translations of the Gaussian induced by the discretization. Conceptually, this can provide a precise intuition for what we term unavoidable canonical nonlinearity within the statistical manifold, which has been in principle invisible effects solely under the KL-divergence based evaluation on a discrete statistical manifold
IV Acknowledgement
[…]

### v2

@@
 We further show that this limiting Wasserstein distance admits a clear geometric interpretation on the statistical manifold, equivalent to a KL divergence associated with the expected parallel translations of continuous Gaussian.
-Our framework thus provides a transport-information-geometric characterization of discretization-induced nonlinearity in classical discrete systems, with future potential applications to configurational thermodynamics.
+Our framework thus provides a transport-information-geometric characterization of discretization-induced nonlinearity in classical discrete systems.
+In addition, we confirm that this $W_{2}$-KL equivalence admits a natural generalization beyond Gaussian families, provided that the transport cost is aligned with the Fisher metric of an underlying statistical submanifold and the discretization scale induces infinitesimal parameter variations.
 I Introduction
@@
 In realistic discrete systems, however, the CDOS inevitably deviates from Gaussian families due to the discrete nature of configuration space imposed by lattice constraints.
-Existing approaches to quantify such canonical nonlinearity are primarily based on information geometry, in particular the Kullback-Leibler (KL) divergence: The nonlinearity was evaluated by comparing the discrete CDOS of a real system with a reference Gaussian distribution having the same mean and covariance matrix, where the Gaussian is discretized on the same configurational support as the real CDOS.
+Existing approaches to quantify such canonical nonlinearity are primarily based on information geometry, in particular the Kullback-Leibler (KL) divergence:klo The nonlinearity was evaluated by comparing the discrete CDOS of a real system with a reference Gaussian distribution having the same mean and covariance matrix, where the Gaussian is discretized on the same configurational support as the real CDOS.
 While this framework captures deviations of the CDOS from Gaussian, it necessarily includes additional contributions arising from the discretization of continuous Gaussian families themselves.
@@
 The present study thus provides a transport-information-geometric characterization of discretization-induced nonlinearity in classical discrete systems, which has remained inaccessible to information-geometric approaches alone.
+In the final part of this paper, we show that the derived $W_{2}$-KL equivalence can be extended to a broader class of statistical submanifolds beyond Gaussian families, under the condition that the transport cost is aligned with the Fisher metric associated with the relevant degrees of freedom in system of interest.
 The details are shown below.
@@
 This choice can be viewed as replacing the standard Euclidean squared distance $\|x-y\|^{2}$ in the definition of the conventional 2-Wasserstein distance by the squared distance induced by the Fisher metric on the Gaussian statistical manifold.
-When we consider Gaussian families, its landscape is completely determined by mean vector $\mu$ and covariance matrix $\Gamma$.
-Considering that (i) $\mu\in\mathbb{R}^{f}$ and $\Gamma\in\mathbf{M}_{f,f}\left(\mathbb{R}\right)$ and (ii) transport object is in $\mathbb{R}^{f}$, appropriate Fisher metric is associated with translations of the mean while fixing covariance matrix, which reduces to the constant metric tensor $\Gamma^{-1}$.
-In fact, under the definition of the cost function of Eq.
-(6), we see the following relationship for Gaussian families:
-$^{\forall}\delta\mu\in\mathbb{R}^{f},\$$W^{2}_{2}\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right)$ (7)
-$=2D\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right),$
-under the restriction of fixed covariance matrix and only allowing translations of its mean.
+Note here that the $W_{2}$ distance between two Gaussian of $g\left(\mu_{1},\Gamma\right)$ and $g\left(\mu_{2},\Gamma\right)$ under the Euclidean metric is known as $\left|\mu_{1}-\mu_{2}\right|^{2}$.we When we consider Gaussian families, its landscape is completely determined by mean vector $\mu$ and covariance matrix $\Gamma$.
+Considering that (i) $\mu\in\mathbb{R}^{f}$ and $\Gamma\in\mathbf{M}_{f,f}\left(\mathbb{R}\right)$ and (ii) transport object is in $\mathbb{R}^{f}$, appropriate Fisher metric is associated with translations of the mean while fixing covariance matrix, which reduces to the constant metric tensor $\Gamma^{-1}$.aig In fact, under the definition of the cost function of Eq.
+(6) with the restriction of fixed covariance matrix and only allowing translations of its mean, $W_{2}^{2}=2D$ holds for any translation, supporting the compatibility of the introduced cost function in Eq.
+(6).
 Under these preparations, we should quantify $W_{2}\left(P_{c},P_{d}\right)$, i.e., 2-Wesserstein distance between a continuous multivariate Gaussian distribution $P_{c}$ with covariance matrix $\Gamma$, and its discretized counterpart $P_{d}$.
@@
 Explicitly, this process is expressed as
-$^{\forall}k,\ P_{d}\left(q_{k}\right)=\int_{V_{k}}P_{c}\left(q\right)dq,\ d\to 0.$ (8)
+$^{\forall}k,\ P_{d}\left(q_{k}\right)=\int_{V_{k}}P_{c}\left(q\right)dq.$ (7)
 In the limit $d\to 0$, the optimal transport plan is expected to converge to a local projection that maps each point $q\in V_{k}$ to the representative point $q^{\prime}_{k}$, namely,
@@
 $\sum_{k}P_{c}\left(q_{k}\right)d^{f}=\int_{\mathbb{R}^{f}}P_{c}\left(q\right)dq=1.$ (11)
-Remarkably, this expression is universal: it depends only on the covariance matrix of the Gaussian distribution and the discretization scale $d$, and is independent of higher-order moments induced by the discretization process.
+Notably, this expression is universal: it depends only on the covariance matrix of the Gaussian distribution and the discretization scale $d$, and is independent of higher-order moments induced by the discretization process.
 This result provides a quantitative measure of the UCN induced purely by discretization.
 Information-Geometric Intepretation of $W_{2}$
-We finally provide an information-geometric interpretation of the derived Wasserstein distance in Eq.
+We then provide an information-geometric interpretation of the derived Wasserstein distance in Eq.
 (10), clarifying its equivalence to the KL divergence associated with expected parallel translations of Gaussian distributions.
@@
 (10), we perform discretization of Gaussian for $f$-dimensional hypercubic with side length $d$.
-At the side of statistical manifold, we first extend the translation magnitude of Gaussian $\delta\mu$ to random vector as an i.i.d.
+At the side of statistical manifold, we first extend the translation magnitude of Gaussian $\delta\mu$ to an i.i.d.
 random vector with probability density $\rho$ given by the uniform distribution, namely,
@@
 $\mathbb{E}_{\rho}\left[\delta\mu^{\textrm{T}}\Gamma^{-1}\delta\mu\right]=\sum_{i,k}\left(\Gamma^{-1}\right)_{ik}\mathbb{E}_{\rho}\left[\delta\mu_{i}\delta\mu_{k}\right].$ (14)
-Since $\mathbb{E}_{\rho}\left[\delta\mu_{i}\delta\mu_{k}\right]$ corresponds to $\left(i,k\right)$-component of the covariance matrix for $\rho\left(\delta\mu\right)$, from Eq. (14),
+Since $\mathbb{E}_{\rho}\left[\delta\mu_{i}\delta\mu_{k}\right]$ corresponds to $\left(i,k\right)$-component of the covariance matrix for $\rho\left(\delta\mu\right)$, from Eq. (13), we can write:
 $\mathbb{E}_{\rho}\left[\delta\mu_{i}\delta\mu_{k}\right]_{i\neq k}=0$
@@
 We therefore obtain the important relationships:
-$\lim_{d\to 0}W_{2}^{2}\left(P_{c}:P_{d}\right)$ $=$ $d^{2}\frac{1}{12}\textrm{Tr}\left(\Gamma^{-1}\right).$
-$=$ $2\cdot E_{\rho}\left[D\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right)\right].$
-Therefore, the squared Wasserstein distance between the continuous Gaussian distribution and its discretized counterpart is equivalent to twice the KL divergence averaged over continuous Gaussian translations generated by the discretization procedure.
+$\lim_{d\to 0}W_{2}^{2}\left(P_{c}:P_{d}\right)$ $=$ $d^{2}\frac{1}{12}\textrm{Tr}\left(\Gamma^{-1}\right)$
+$=$ $2\cdot\mathbb{E}_{\rho}\left[D\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right)\right].$
+Eq.(II) certainly exhibits that the squared Wasserstein distance between the continuous Gaussian distribution and its discretized counterpart is equivalent to twice the KL divergence averaged over continuous Gaussian translations generated by the discretization procedure.
 This result establishes a clear geometric interpretation of the UCN: The cost to project continuous Gaussian onto a discrete configurational support is given by the cumulative information-geometric distance for expected parallel translations within the Gaussian statistical manifold.
-These contributions are invisible to KL-based comparisons restricted to discrete statistical manifolds, but are naturally captured by the present transport-information-geometric framework.
+These contributions are invisible to existing KL-based comparisons restricted to discrete statistical manifolds, but are naturally captured by the present transport-information-geometric framework.
+Generalization of the $W_{2}$-KL Equivalence beyond Gaussian Families
+In the previous sections, we demonstrated that the discrepancy between continuous Gaussian distribution and its discretized counterpart, measured by the 2-Wasserstein distance with a quadratic cost aligned with Fisher metric, converges in the vanishing discretization limit $d\to 0$ to the expectation of the KL divergence between infinitesimally translated continous Gaussian.
+Although these derivations was shown for Gaussian families, the underlying mathematical structure indicates its broader applicability.
+We here briefly address the conditions where such a correspondence can be generalized.
+Let $S$ be a statistical submanifold for a certain set of continous probability functions, (wholely or partly) parameterized by coordinates of
+$\xi=\left(\xi_{1},\cdots,\xi_{f}\right),$ (18)
+where $f$ corresponds to the number of effective degrees of freedom relevant to the description of the system interested (and, subsequently, to the transport problem).
+Then we consider the problem to measure the difference between function $P\in S$ and its discretized counterpart $P^{\prime}$, based on $W_{2}$ and KL-divergence as discussed for Gaussian families.
+We assume that, in these coordinates, the Fisher metric (or its block-diagonal component associated with the transported degrees of freedom) is given by $f\times f$ positive-definite matrix of $\Omega$, which is taken as constant.
+More generally, our statement rely on the locally-constant character of $\Omega$ under inifitesimal variations in $\xi$.
+Then we define the quadratic cost function for the 2-Wasserstein distance $W_{2}\left(P:P^{\prime}\right)$ as
+$c\left(q,q^{\prime}\right)=\left(q-q^{\prime}\right)^{\textrm{T}}\Omega\left(q-q^{\prime}\right),$ (19)
+which ensures consistency between the transport cost and the local information-geometric structure of $S$.
+Then, we discretize the underlying continuous space for $P$, by using $f$-dimensional hypercubic with side length $d$, and consider its limit of $d\to 0$.
+At the same time, we assume that the uncertainty or variation of the parameters $\xi$, induced by the discretization scales $d$, can be given by
+$\delta\xi=O\left(h\left(d\right)\right),\ h\left(d\right)=a\cdot d,$ (20)
+where $a>0$ denotes a scalar constant, typically characterizing the sensitivity of the parameters $\xi$ to the discretization scale.
+Then we interpreted $\delta\xi$ as the following i.i.d.
+random vector with its distribution taking $\rho$:
+$\rho\left(\delta\xi\right)=\textrm{Unif}\left[-\frac{h\left(d\right)}{2},\frac{h\left(d\right)}{2}\right]^{f}\ \left(\textrm{i.i.d.}\right).$ (21)
+Note that the form of the distribution is not necessarily restricted to uniform one, where its variances scale as $O\left(d^{2}\right)$ (i.i.d.) is essential: This only affects the following $W_{2}$-KL equivalence up to the constant factor.
+Under these preparations, let us recall that KL-divergence between two nearby distributions in $S$ admits the following expansion:
+$D\left(P_{\xi+\delta\xi}:P_{\xi}\right)=\frac{1}{2}\delta\xi^{\textrm{T}}\Omega\delta\xi+o\left(\left\|\delta\xi\right\|^{2}\right),$ (22)
+where in the case of Gaussian families with $\xi=\mu$, the r.h.s.
+is exactly given by up to the quadratic form.
+Now it is clear that when we take expectation of KL-divergence in Eq.
+(22) w.r.t.
+$\rho\left(\delta\xi\right)$, its leading order at $d\to 0$ takes
+$\mathbb{E}_{\rho\left(\delta\xi\right)}\left[D\left(P_{\xi+\delta\xi}:P_{\xi}\right)\right]=\frac{a^{2}d^{2}}{24}\textrm{Tr}\left(\Omega\right).$ (23)
+For the side of optimal transport theory, $W_{2}$ distance can be obtained in the same way as Gaussian families, namely,
+$\lim_{d\to 0}W_{2}^{2}=\frac{d^{2}}{12}\textrm{Tr}\left(\Omega\right).$ (24)
+Thereby, we obtain a generalized $W_{2}$-KL equivalence within the leading order at $d\to 0$:
+$W_{2}^{2}\left(P:P^{\prime}\right)\sim\frac{2}{a^{2}}\mathbb{E}_{\rho\left(\delta\xi\right)}\left[D\left(P_{\xi+\delta\xi}:P_{\xi}\right)\right].$ (25)
+The proportionality constant explicitly reflects the scaling relation $h\left(d\right)=a\cdot d$, between the discretization length and the induced parameter uncertainty or its variation.
+For Gaussian families, where the coordinates $\xi$ corresponds to the mean vector $\mu$, and $a=1$, this equivalence provides a direct physical intuition in terms of the Gaussian translations as discussed.
+In more general statistical submanifolds, the same structure should be interpreted as an information-geometric statement: the discretization-induced cost measured by 2-Wasserstein distance extracts the second-order geometric structure of the KL divergence in terms of Fisher metric at the leading order of $d\to 0$, which is independent of the detailed functional form of the distributions.
 III Conclusions
@@
 Our central result exhibits that the derived $W_{2}$ distance admits a clear information-geometric interpretation: it is exactly equivalent to twice the Kullback-Leibler divergence averaged over random parallel translations of the Gaussian induced by the discretization.
-Conceptually, this can provide a precise intuition for what we term unavoidable canonical nonlinearity within the statistical manifold, which has been in principle invisible effects solely under the KL-divergence based evaluation on a discrete statistical manifold
+Conceptually, this can provide a precise intuition for what we term unavoidable canonical nonlinearity within the statistical manifold, which has been in principle invisible effects solely under the KL-divergence based evaluation on a discrete statistical manifold.
+Finally, the present results indicate that the leading-order correspondence between the discretization-induced 2-Wasserstein distance and the expectation of the KL divergence reflects a general transport-information-geometric structure associated with the statistical submanifolds of interest.
+The unavoidable canonical nonlinearity observed in continuous Gaussian families and their discretized counterparts should therefore be understood as a particularly transparent example of this more general correspondence, where the parameter variations admit a clear physical interpretation in terms of parallel translations of Gaussian distributions.
 IV Acknowledgement

### v3

@@
 These contributions are invisible to existing KL-based comparisons restricted to discrete statistical manifolds, but are naturally captured by the present transport-information-geometric framework.
+In a realistic discrete system on a lattice, the following three features are unavoidable: (i) $d$ takes a nonzero positive value, (ii) the domain of the CDOS is bounded, and (iii) the domain is generally asymmetric.
+All of these features are dominated by the underlying lattice.
+Therefore, to address the effect of UCN in realistic systems, numerical approaches, such as systematic comparisons between Eq.
+(II) and Eq.
+(8) under various conditions, are essential for disentangling the individual contributions of these three lattice-induced effects.
 Generalization of the $W_{2}$-KL Equivalence beyond Gaussian Families
-In the previous sections, we demonstrated that the discrepancy between continuous Gaussian distribution and its discretized counterpart, measured by the 2-Wasserstein distance with a quadratic cost aligned with Fisher metric, converges in the vanishing discretization limit $d\to 0$ to the expectation of the KL divergence between infinitesimally translated continous Gaussian.
+In the previous sections, we demonstrated that the discrepancy between continuous Gaussian distribution and its discretized counterpart, measured by the 2-Wasserstein distance with a quadratic cost aligned with Fisher metric, converges in the vanishing discretization limit $d\to 0$ to the expectation of the KL divergence between infinitesimally translated continuous Gaussian.
 Although these derivations was shown for Gaussian families, the underlying mathematical structure indicates its broader applicability.
@@
 We assume that, in these coordinates, the Fisher metric (or its block-diagonal component associated with the transported degrees of freedom) is given by $f\times f$ positive-definite matrix of $\Omega$, which is taken as constant.
-More generally, our statement rely on the locally-constant character of $\Omega$ under inifitesimal variations in $\xi$.
+More generally, our statement rely on the locally-constant character of $\Omega$ under infinitesimal variations in $\xi$.
 Then we define the quadratic cost function for the 2-Wasserstein distance $W_{2}\left(P:P^{\prime}\right)$ as
@@
 random vector with its distribution taking $\rho$:
-$\rho\left(\delta\xi\right)=\textrm{Unif}\left[-\frac{h\left(d\right)}{2},\frac{h\left(d\right)}{2}\right]^{f}\ \left(\textrm{i.i.d.}\right).$ (21)
-Note that the form of the distribution is not necessarily restricted to uniform one, where its variances scale as $O\left(d^{2}\right)$ (i.i.d.) is essential: This only affects the following $W_{2}$-KL equivalence up to the constant factor.
+$\rho\left(\delta\xi\right)=\textrm{Unif}\left[-\frac{h\left(d\right)}{2},\frac{h\left(d\right)}{2}\right]^{f}\ \left(\textrm{i.i.d.}\right),$ (21)
+where the form of the distribution is not necessarily restricted to uniform one, where its variances scale as $O\left(d^{2}\right)$ (i.i.d.) is essential: This only affects the following $W_{2}$-KL equivalence up to the constant factor.
 Under these preparations, let us recall that KL-divergence between two nearby distributions in $S$ admits the following expansion:
@@
 Thereby, we obtain a generalized $W_{2}$-KL equivalence within the leading order at $d\to 0$:
-$W_{2}^{2}\left(P:P^{\prime}\right)\sim\frac{2}{a^{2}}\mathbb{E}_{\rho\left(\delta\xi\right)}\left[D\left(P_{\xi+\delta\xi}:P_{\xi}\right)\right].$ (25)
-The proportionality constant explicitly reflects the scaling relation $h\left(d\right)=a\cdot d$, between the discretization length and the induced parameter uncertainty or its variation.
-For Gaussian families, where the coordinates $\xi$ corresponds to the mean vector $\mu$, and $a=1$, this equivalence provides a direct physical intuition in terms of the Gaussian translations as discussed.
+$W_{2}^{2}\left(P:P^{\prime}\right)=\frac{2}{a^{2}}\mathbb{E}_{\rho\left(\delta\xi\right)}\left[D\left(P_{\xi+\delta\xi}:P_{\xi}\right)\right].$ (25)
+Remarks on the Generalization
+The proportionality constant in Eq.
+(25) explicitly reflects the scaling relation $h\left(d\right)=a\cdot d$, between the discretization length and the induced parameter uncertainty or its variation.
+For Gaussian families, where the coordinates $\xi$ corresponds to the mean vector $\mu$ and $a=1$, this equivalence provides a direct physical intuition in terms of the Gaussian translations as discussed.
 In more general statistical submanifolds, the same structure should be interpreted as an information-geometric statement: the discretization-induced cost measured by 2-Wasserstein distance extracts the second-order geometric structure of the KL divergence in terms of Fisher metric at the leading order of $d\to 0$, which is independent of the detailed functional form of the distributions.
+We emphasize that the above leading-order correspondence relies on the assumption that the discretization-induced transport cost is dominated by local fluctuations: This condition can be naturally satisfied for distributions with sufficiently fast-decaying tails (including Gaussian families), where distant contributions to the quadratic transport cost are reasonably suppressed.
+In contrast, for distributions exhibiting heavy tails or slow decays, non-local contributions from the tail region may modify the scaling behavior of the $W_{2}$ distance in the limit $d\to 0$, potentially requiring corrections including the higher-order terms in $d$.
+The present analysis therefore characterizes a universal local correspondence, whose extension to heavy-tailed distributions is left for our future work.
+Through the discussions about generalization, we now see that the extension of $\delta\xi$ to random vector is not merely a matter of average, but is structurally necessary: The trace structure appearing in the $W_{2}$ of Eq.
+(24) can only emerge in quadratic form of the KL divergence, if the outer product of $\delta\xi$ is effectively proportional to the identity matrix $I$, i.e., $\delta\xi\delta\xi^{\mathsf{T}}\propto d^{2}I_{f}$.
+Since the outer product of a single vector is necessarily rank one, while the required structure is full rank with $f$ degrees of freedom, an extension to a probabilistic ensemble of $\delta\xi$ is unavoidable for systems with $f\geq 2$.
+Finally, we comment on the role of parametrization and invariance in the present Wasserstein-KL correspondence.
+While the KL divergence itself is invariant under reparametrizations on the statistical manifold, Eqs.
+(24)-(25) suggest that its expectation under $\rho$ depends on the trace of Fisher metric $\Omega$.
+This does not represent a contradiction, but can be reasonably understood from the following two aspects.
+(i) Taking the expectation $\mathbb{E}_{\rho}$ in Eq.
+(25) is an extrinsic operation living outside the statistical manifold, where the infinitesimal change in the selected parameter, $\delta\xi$, is linked to the discretization scale $d$ and subsequently extended to a random vector (the associated probability density $\rho$ typically reflects a modeling choice related to discretization or coarse-graining, depending on the problem under consideration).
+(ii) In general, when higher-order terms are present, the KL divergence expanded up to second order in $\delta\xi$, yielding a quadratic form, is explicitly governed by the Fisher metric $\Omega$ associated with the chosen parameter $\xi$.
+In such cases, once combined with the discretization scale and the extrinsic expectation over parameter variations, the resulting numerical value depends on the chosen parameter and Fisher metric.
+From these viewpoints, the explicit expression $d^{2}\mathrm{Tr}(\Omega)/12$ generally depends on the chosen parameter, yet the correspondence between the 2-Wasserstein transport cost and the expected KL divergence should be understood as an invariant structural relation.
+In particular, for the case of UCN in this work, since the Fisher metric is naturally identified through the isomorphism between the parameter $\mu$ on the statistical manifold and the transport coordinate, and the parameter variations scale linearly with the discretization length, the proportionality constant can be fixed.
+As a result, the correspondence yields an equality, Eq.
+(II), rather than a mere proportionality.
+In more general settings, different parameters and modeling choices for parameter variations lead to different numerical prefactors.
+Nevertheless, the universal correspondence $W_{2}^{2}\propto 2/a^{2}\cdot\mathbb{E}_{\rho}[D]$ is always preserved, expressing the robust emergence of second-order information-geometric structure from discretization-induced transport costs.
 III Conclusions

### v4

@@
 Our framework thus provides a transport-information-geometric characterization of discretization-induced nonlinearity in classical discrete systems.
-In addition, we confirm that this $W_{2}$-KL equivalence admits a natural generalization beyond Gaussian families, provided that the transport cost is aligned with the Fisher metric of an underlying statistical submanifold and the discretization scale induces infinitesimal parameter variations.
+In addition, we confirm that this $W_{2}$-KL equivalence admits a natural generalization beyond Gaussian families, provided that the transport cost is aligned with the Fisher metric of an underlying statistical submanifold and the discretization scale links to infinitesimal parameter variations.
 I Introduction
@@
 Such discretization-induced effects remain invisible on discrete statistical manifolds, where probability measures are compared only on a fixed set of discrete supports.
-In the present work, we identify this contribution as an unavoidable canonical nonlinearity (UCN).
+This limitation ultimately originates from the absence of a framework that allows one to interpret discretization-induced geometric distortions of continuous distributions within information geometry.
+In the present work, we define the discretization-induced geometric contribution for Gausian families as an unavoidable canonical nonlinearity (UCN), and demonstrate that it admits a consistent information-geometric interpretation, thereby addressing this missing piece.
 To quantify the UCN, a framework is required that is sensitive to probability weights as well as to the geometric rearrangement of probability mass induced by discretization.
@@
 Since the KL divergence requires the two distributions to be defined on the same support (or to satisfy an inclusion relationship of supports), the continuous Gaussian $P_{c}$ has been discretized on the same configurational support as $P$, yielding a discretized Gaussian distribution $P_{d}$.
-Accordingly, canonical nonlinearity has been characterized by the KL divergence $D\left(P:P_{d}\right)$, which extends to the KL divergence between the canonical distributions induced by the CDOSs $P$ and $P_{d}$.
+Accordingly, canonical nonlinearity has been characterized by the KL divergence $D\left(P:P_{d}\right)$, which subsequently extends to the KL divergence between the canonical distributions induced by the CDOSs $P$ and $P_{d}$.
 While this framework successfully captures non-Gaussian features inherent to the realistic CDOS, it inevitably includes an additional contribution originating from the discretization of the continuous Gaussian distribution itself.
@@
 (6).
-Under these preparations, we should quantify $W_{2}\left(P_{c},P_{d}\right)$, i.e., 2-Wesserstein distance between a continuous multivariate Gaussian distribution $P_{c}$ with covariance matrix $\Gamma$, and its discretized counterpart $P_{d}$.
+Under these preparations, we should quantify $W_{2}\left(P_{c},P_{d}\right)$, i.e., 2-Wesserstein distance between a continuous multivariate Gaussian distribution $P_{c}$ with mean vector $\mu$ and covariance matrix $\Gamma$, and its discretized counterpart $P_{d}$.
 Here, we consider discretization process as partitioning the configuration space ($\simeq\mathbb{R}^{f}$) into hypercubic cells $V_{k}$ of side length $d$, then taking the vanishing discretization limit of $d\to 0$.
@@
 We therefore obtain the important relationships:
-$\lim_{d\to 0}W_{2}^{2}\left(P_{c}:P_{d}\right)$ $=$ $d^{2}\frac{1}{12}\textrm{Tr}\left(\Gamma^{-1}\right)$
-$=$ $2\cdot\mathbb{E}_{\rho}\left[D\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right)\right].$
-Eq.(II) certainly exhibits that the squared Wasserstein distance between the continuous Gaussian distribution and its discretized counterpart is equivalent to twice the KL divergence averaged over continuous Gaussian translations generated by the discretization procedure.
+$\lim_{d\to 0}W_{2}^{2}\left(P_{c}\left(\mu,\Gamma\right):P_{d}\right)$ $=$ $2\cdot\mathbb{E}_{\rho}\left[D\left(P_{c}\left(\mu+\delta\mu,\Gamma\right):P_{c}\left(\mu,\Gamma\right)\right)\right]$ (17)
+$=$ $d^{2}\frac{1}{12}\textrm{Tr}\left(\Gamma^{-1}\right).\quad$
+Eq.(17) certainly exhibits that the squared Wasserstein distance between the continuous Gaussian distribution and its discretized counterpart is equivalent to twice the KL divergence averaged over continuous Gaussian translations generated by the discretization procedure.
 This result establishes a clear geometric interpretation of the UCN: The cost to project continuous Gaussian onto a discrete configurational support is given by the cumulative information-geometric distance for expected parallel translations within the Gaussian statistical manifold.
@@
 Therefore, to address the effect of UCN in realistic systems, numerical approaches, such as systematic comparisons between Eq.
-(II) and Eq.
+(17) and Eq.
 (8) under various conditions, are essential for disentangling the individual contributions of these three lattice-induced effects.
@@
 where $f$ corresponds to the number of effective degrees of freedom relevant to the description of the system interested (and, subsequently, to the transport problem).
-Then we consider the problem to measure the difference between function $P\in S$ and its discretized counterpart $P^{\prime}$, based on $W_{2}$ and KL-divergence as discussed for Gaussian families.
+Then we consider the problem to measure the difference between function $P_{\xi}\in S$ and its discretized counterpart $P^{\prime}$, based on $W_{2}$ and KL-divergence as discussed for Gaussian families.
 We assume that, in these coordinates, the Fisher metric (or its block-diagonal component associated with the transported degrees of freedom) is given by $f\times f$ positive-definite matrix of $\Omega$, which is taken as constant.
 More generally, our statement rely on the locally-constant character of $\Omega$ under infinitesimal variations in $\xi$.
-Then we define the quadratic cost function for the 2-Wasserstein distance $W_{2}\left(P:P^{\prime}\right)$ as
+Now we define the quadratic cost function for the 2-Wasserstein distance $W_{2}\left(P_{\xi}:P^{\prime}\right)$ as
 $c\left(q,q^{\prime}\right)=\left(q-q^{\prime}\right)^{\textrm{T}}\Omega\left(q-q^{\prime}\right),$ (19)
@@
 Then, we discretize the underlying continuous space for $P$, by using $f$-dimensional hypercubic with side length $d$, and consider its limit of $d\to 0$.
-At the same time, we assume that the uncertainty or variation of the parameters $\xi$, induced by the discretization scales $d$, can be given by
+Under this discretization, we introduce the following linear relationship betwen $\delta\xi$ and $d$:
 $\delta\xi=O\left(h\left(d\right)\right),\ h\left(d\right)=a\cdot d,$ (20)
-where $a>0$ denotes a scalar constant, typically characterizing the sensitivity of the parameters $\xi$ to the discretization scale.
-Then we interpreted $\delta\xi$ as the following i.i.d.
+where $a>0$ reflects the relative scaling arising from the identification of the local quadratic structure of the KL divergence with the discretization-induced transport cost.
+Then we extend $\delta\xi$ to the following i.i.d.
 random vector with its distribution taking $\rho$:
@@
 Thereby, we obtain a generalized $W_{2}$-KL equivalence within the leading order at $d\to 0$:
-$W_{2}^{2}\left(P:P^{\prime}\right)=\frac{2}{a^{2}}\mathbb{E}_{\rho\left(\delta\xi\right)}\left[D\left(P_{\xi+\delta\xi}:P_{\xi}\right)\right].$ (25)
+$W_{2}^{2}\left(P_{\xi}:P^{\prime}\right)=\frac{2}{a^{2}}\mathbb{E}_{\rho\left(\delta\xi\right)}\left[D\left(P_{\xi+\delta\xi}:P_{\xi}\right)\right].$ (25)
 Remarks on the Generalization
-The proportionality constant in Eq.
-(25) explicitly reflects the scaling relation $h\left(d\right)=a\cdot d$, between the discretization length and the induced parameter uncertainty or its variation.
-For Gaussian families, where the coordinates $\xi$ corresponds to the mean vector $\mu$ and $a=1$, this equivalence provides a direct physical intuition in terms of the Gaussian translations as discussed.
-In more general statistical submanifolds, the same structure should be interpreted as an information-geometric statement: the discretization-induced cost measured by 2-Wasserstein distance extracts the second-order geometric structure of the KL divergence in terms of Fisher metric at the leading order of $d\to 0$, which is independent of the detailed functional form of the distributions.
+While the W2-KL correspondence in Eq.
+(17) provides a transparent physical intuition, i.e., the transport cost between continuous and discrete Gaussian is interpreted as the expected parallel translation of continuous Gaussian, in more general statistical submanifolds, the same structure should be interpreted as an information-geometric statement.
+The discretization-induced cost measured by 2-Wasserstein distance extracts the second-order geometric structure of the KL divergence in terms of Fisher metric at the leading order of $d\to 0$, which is independent of the detailed functional form of the distributions.
 We emphasize that the above leading-order correspondence relies on the assumption that the discretization-induced transport cost is dominated by local fluctuations: This condition can be naturally satisfied for distributions with sufficiently fast-decaying tails (including Gaussian families), where distant contributions to the quadratic transport cost are reasonably suppressed.
@@
 The present analysis therefore characterizes a universal local correspondence, whose extension to heavy-tailed distributions is left for our future work.
-Through the discussions about generalization, we now see that the extension of $\delta\xi$ to random vector is not merely a matter of average, but is structurally necessary: The trace structure appearing in the $W_{2}$ of Eq.
+Through the discussions about Gaussian families and its generalization, we now see that the extension of $\delta\xi$ to random vector is not merely a matter of average, but is structurally necessary: The trace structure appearing in the $W_{2}$ of Eq.
 (24) can only emerge in quadratic form of the KL divergence, if the outer product of $\delta\xi$ is effectively proportional to the identity matrix $I$, i.e., $\delta\xi\delta\xi^{\mathsf{T}}\propto d^{2}I_{f}$.
 Since the outer product of a single vector is necessarily rank one, while the required structure is full rank with $f$ degrees of freedom, an extension to a probabilistic ensemble of $\delta\xi$ is unavoidable for systems with $f\geq 2$.
-Finally, we comment on the role of parametrization and invariance in the present Wasserstein-KL correspondence.
+We also comment on the role of parametrization and invariance in the present Wasserstein-KL correspondence.
 While the KL divergence itself is invariant under reparametrizations on the statistical manifold, Eqs.
@@
 From these viewpoints, the explicit expression $d^{2}\mathrm{Tr}(\Omega)/12$ generally depends on the chosen parameter, yet the correspondence between the 2-Wasserstein transport cost and the expected KL divergence should be understood as an invariant structural relation.
-In particular, for the case of UCN in this work, since the Fisher metric is naturally identified through the isomorphism between the parameter $\mu$ on the statistical manifold and the transport coordinate, and the parameter variations scale linearly with the discretization length, the proportionality constant can be fixed.
-As a result, the correspondence yields an equality, Eq.
-(II), rather than a mere proportionality.
-In more general settings, different parameters and modeling choices for parameter variations lead to different numerical prefactors.
-Nevertheless, the universal correspondence $W_{2}^{2}\propto 2/a^{2}\cdot\mathbb{E}_{\rho}[D]$ is always preserved, expressing the robust emergence of second-order information-geometric structure from discretization-induced transport costs.
+The exception is the Gaussian families considered in the UCN analysis, where (i) the Fisher metric is naturally identified through the isomorphism between the parameter $\mu$ on the statistical manifold and the transport coordinate, and (ii) the natural parameterization and the standard quadratic transport cost lead to $a=1$.
+While this value of $a$ is not mathematically unique, it corresponds to a canonical choice in which KL divergence on the statistical manifold and the transport cost induce the same local quadratic form, so that no additional rescaling is required.
+In more general settings, different choices for parameters typically lead to different numerical prefactors.
+Nevertheless, the universal correspondence $W_{2}^{2}\propto 2\cdot\mathbb{E}_{\rho}[D]$ is always preserved, expressing the robust emergence of second-order information-geometric structure from discretization-induced transport costs.
+The explicit prefactor depends only on the extrinsic identification between the discretization scale and infinitesimal parameter variations.
+Conceptual Position
+In light of the above remarks, it is instructive to clarify the conceptual position of the present approach.
+The central aim of this work is to clarify how discretization-induced nonlinearity (UCN), inherent to classical discrete systems, can be meaningfully characterized from an information-geometric perspective.
+At first sight, this objective may appear ill-posed: a continuous probability distribution and its discretized counterpart generally belong to different measure classes and possess incompatible supports, so that their discrepancy cannot be directly quantified within a single statistical manifold using conventional divergences such as the KL divergence.
+In this sense, a direct information-geometric comparison between continuous and discrete distributions is fundamentally obstructed.
+To overcome this difficulty, we do not attempt to directly measure the discrepancy between continuous and discretized distributions themselves.
+Instead, we relax the original problem and focus on the geometric distortion induced by the discretization procedure viewed as an extrinsic projection from a continuous configuration space onto a discrete lattice.
+While this projection operation is not definable intrinsically on a statistical manifold, its effect can nevertheless be quantified by introducing an appropriate transport metric on the underlying sample space.
+The key idea is then to interpret the resulting transport cost by pulling it back to the statistical manifold through infinitesimal parameter variations.
+Under discretization, the lattice scale $d$ and infinitesimal parameter variations $\delta\xi$ capture a common notion of scale dependence: On the statistical manifold, $\delta\xi$ controls the distinguishability of nearby distributions, while on the level of the underlying sample space, $d$ governs the transport cost that measures geometric distortions between their supports.
+To consistently quantify this scale-dependent correspondence on the transport side, its cost function should encode the geometric information controlling the infinitesimal distinguishability on the statistical manifold, i.e., employing a quadratic cost canonically consistent with the local Fisher metric.
+From this correspondence, $d$ is not merely a numerical resolution parameter, but is explicitly linked to $\delta\xi$.
+The resultant 2-Wasserstein distance naturally provides a differential, second-order characterization of the distortion induced by the projection, admits a well-defined limit as $d\to 0$, and allows a direct information-geometric interpretation.
+The $W_{2}$-KL equivalence should therefore be understood not as an artificial identification of unrelated quantities, but as a consistent pullback of an extrinsic geometric operation into an intrinsic, Fisher-metric-based description on the statistical manifold.
+In this sense, the present framework offers a natural and minimal route to interpret discretization-induced nonlinearity in terms of information geometry.
+Possible alternative metrics or higher-order corrections may encode additional geometric structures; nevertheless, under the requirements of differentiability, quadratic approximability, and a controlled discretization limit, the present transport-information-geometric construction emerges as one of the most natural choices.
 III Conclusions

### v5

@@
 This limitation ultimately originates from the absence of a framework that allows one to interpret discretization-induced geometric distortions of continuous distributions within information geometry.
-In the present work, we define the discretization-induced geometric contribution for Gausian families as an unavoidable canonical nonlinearity (UCN), and demonstrate that it admits a consistent information-geometric interpretation, thereby addressing this missing piece.
+In the present work, we define the discretization-induced geometric contribution for Gaussian families as an unavoidable canonical nonlinearity (UCN), and demonstrate that it admits a consistent information-geometric interpretation, thereby addressing this missing piece.
 To quantify the UCN, a framework is required that is sensitive to probability weights as well as to the geometric rearrangement of probability mass induced by discretization.
@@
 The present study thus provides a transport-information-geometric characterization of discretization-induced nonlinearity in classical discrete systems, which has remained inaccessible to information-geometric approaches alone.
-In the final part of this paper, we show that the derived $W_{2}$-KL equivalence can be extended to a broader class of statistical submanifolds beyond Gaussian families, under the condition that the transport cost is aligned with the Fisher metric associated with the relevant degrees of freedom in system of interest.
-The details are shown below.
-II Concept and Derivation
-Unavoidable Canonical Nonliearity
+In the following part of this paper, we show that the derived $W_{2}$-KL equivalence can be extended to a broader class of statistical submanifolds beyond Gaussian families, under the condition that the transport cost is aligned with the Fisher metric associated with the relevant degrees of freedom in system of interest.
+We will see that while the present mathematical derivations are elementary, the underlying conceptual structure relies on nontrivial identifications; readers primarily interested in the interpretation may wish to consult the Conceptual Position in the final part.
+II Derivation and Concept
+Unavoidable Canonical Nonlinearity
 Here, we briefly clarify the concept of canonical nonlinearity and its unavoidable contribution, referred to as the unavoidable canonical nonlinearity (UCN).
@@
 These considerations naturally motivate the present use of optimal transport theory.
-Wesserstein Distance Aligned with Fisher Metric
+Wasserstein Distance Aligned with Fisher Metric
 Optimal transport theory can provide a natural framework to compare probability distributions defined on different supports, e.g., continuous and discrete ones.
@@
 (6).
-Under these preparations, we should quantify $W_{2}\left(P_{c},P_{d}\right)$, i.e., 2-Wesserstein distance between a continuous multivariate Gaussian distribution $P_{c}$ with mean vector $\mu$ and covariance matrix $\Gamma$, and its discretized counterpart $P_{d}$.
+Under these preparations, we should quantify $W_{2}\left(P_{c},P_{d}\right)$, i.e., 2-Wasserstein distance between a continuous multivariate Gaussian distribution $P_{c}$ with mean vector $\mu$ and covariance matrix $\Gamma$, and its discretized counterpart $P_{d}$.
 Here, we consider discretization process as partitioning the configuration space ($\simeq\mathbb{R}^{f}$) into hypercubic cells $V_{k}$ of side length $d$, then taking the vanishing discretization limit of $d\to 0$.
 Explicitly, this process is expressed as
-$^{\forall}k,\ P_{d}\left(q_{k}\right)=\int_{V_{k}}P_{c}\left(q\right)dq.$ (7)
+$^{\forall}k,\ P_{d}\left(q^{\prime}_{k}\right)=\int_{V_{k}}P_{c}\left(q\right)dq.$ (7)
 In the limit $d\to 0$, the optimal transport plan is expected to converge to a local projection that maps each point $q\in V_{k}$ to the representative point $q^{\prime}_{k}$, namely,
@@
 $u=q-q^{\prime}_{k},\ ^{\forall}i:\ \left|u_{i}\right|\leq\frac{d}{2}$ (9)
-with representative hypercubic as $V_{0}$, and taking Tailor series expantion of $P_{c}\left(q^{\prime}_{k}+u\right)$ around $q^{\prime}_{k}$, we can reasonably retain the leading order in $d$, thereby
+with representative hypercubic as $V_{0}$, and taking Taylor series expantion of $P_{c}\left(q^{\prime}_{k}+u\right)$ around $q^{\prime}_{k}$, we can reasonably retain the leading order in $d$, thereby
 $W_{2}^{2}\left(P_{c}:P_{d}\right)$ $=$ $\sum_{k}P_{c}\left(q_{k}\right)\int_{V_{0}}u^{\textrm{T}}\Gamma^{-1}udu$ (10)
@@
 $\sum_{k}P_{c}\left(q_{k}\right)d^{f}=\int_{\mathbb{R}^{f}}P_{c}\left(q\right)dq=1.$ (11)
+A brief numerical check of the derived expression of Eq. (10) against the original integral form of Eq. (8) is provided in Appendix.
 Notably, this expression is universal: it depends only on the covariance matrix of the Gaussian distribution and the discretization scale $d$, and is independent of higher-order moments induced by the discretization process.
@@
 Under this discretization, we introduce the following linear relationship betwen $\delta\xi$ and $d$:
-$\delta\xi=O\left(h\left(d\right)\right),\ h\left(d\right)=a\cdot d,$ (20)
+$^{\forall}k:\ \delta\xi_{k}=O\left(h\left(d\right)\right),\ h\left(d\right)=a\cdot d,$ (20)
 where $a>0$ reflects the relative scaling arising from the identification of the local quadratic structure of the KL divergence with the discretization-induced transport cost.
@@
 (i) Taking the expectation $\mathbb{E}_{\rho}$ in Eq.
-(25) is an extrinsic operation living outside the statistical manifold, where the infinitesimal change in the selected parameter, $\delta\xi$, is linked to the discretization scale $d$ and subsequently extended to a random vector (the associated probability density $\rho$ typically reflects a modeling choice related to discretization or coarse-graining, depending on the problem under consideration).
+(25) is an extrinsic operation living outside the statistical manifold, where the infinitesimal change in the selected parameter, $\delta\xi$, is linked to the discretization scale $d$ and subsequently extended to a random vector.
 (ii) In general, when higher-order terms are present, the KL divergence expanded up to second order in $\delta\xi$, yielding a quadratic form, is explicitly governed by the Fisher metric $\Omega$ associated with the chosen parameter $\xi$.
@@
 The key idea is then to interpret the resulting transport cost by pulling it back to the statistical manifold through infinitesimal parameter variations.
-Under discretization, the lattice scale $d$ and infinitesimal parameter variations $\delta\xi$ capture a common notion of scale dependence: On the statistical manifold, $\delta\xi$ controls the distinguishability of nearby distributions, while on the level of the underlying sample space, $d$ governs the transport cost that measures geometric distortions between their supports.
-To consistently quantify this scale-dependent correspondence on the transport side, its cost function should encode the geometric information controlling the infinitesimal distinguishability on the statistical manifold, i.e., employing a quadratic cost canonically consistent with the local Fisher metric.
-From this correspondence, $d$ is not merely a numerical resolution parameter, but is explicitly linked to $\delta\xi$.
+Under discretization, the lattice scale $d$ and infinitesimal parameter variations $\delta\xi$ capture a common notion of scale dependence.
+On the statistical manifold, $\delta\xi$ controls the distinguishability of nearby distributions, while on the level of the underlying sample space, $d$ governs the transport cost measuring geometric distortions between their supports.
+To consistently quantify this scale-dependent correspondence on the transport side, the cost function should encode the geometric information that controls infinitesimal distinguishability on the statistical manifold, namely by employing a quadratic cost canonically consistent with the local Fisher metric.
+From this correspondence, $d$ is not merely a discretization parameter, but is explicitly linked to $\delta\xi$.
+Although the link between $d$ and $\delta\xi$, and the subsequent extension of $\delta\xi$ to a random vector can be merged into a single operation, we deliberately separate them to clarify their conceptual roles.
+Within the present transport-information-geometric framework, the probability density $\rho\left(\delta\xi\right)$ is intrinsically constrained only through its second moment, whose scaling $O\left(d^{2}\right)$ encodes the common discretization scale shared by statistical distinguishability and transport cost.
+Under this constraint, the detailed functional form of $\rho\left(\delta\xi\right)$ remains underdetermined.
+This residual freedom reflects the fact that higher-order structures are not fixed by the intrinsic geometry alone.
+Instead, they may be selected once additional, extrinsic physical requirements—such as the microscopic origin of the underlying continuous distributions—are specified.
 The resultant 2-Wasserstein distance naturally provides a differential, second-order characterization of the distortion induced by the projection, admits a well-defined limit as $d\to 0$, and allows a direct information-geometric interpretation.
 The $W_{2}$-KL equivalence should therefore be understood not as an artificial identification of unrelated quantities, but as a consistent pullback of an extrinsic geometric operation into an intrinsic, Fisher-metric-based description on the statistical manifold.
-In this sense, the present framework offers a natural and minimal route to interpret discretization-induced nonlinearity in terms of information geometry.
+In these senses, the present framework offers a natural and minimal route to interpret discretization-induced nonlinearity in terms of information geometry.
 Possible alternative metrics or higher-order corrections may encode additional geometric structures; nevertheless, under the requirements of differentiability, quadratic approximability, and a controlled discretization limit, the present transport-information-geometric construction emerges as one of the most natural choices.
@@
 […]
+Appendix: Numerical verification of the analytic expression for $W_{2}$ at finite $d$
+Figure 1: Left: $d$-dependence of the squared 2-Wasserstein distance $W_{2}^{2}$ obtained from the analytic expression Eq.
+(10) (“Analytic”) and from the numerical evaluation of Eq.
+(8) (“Numerical”).
+Right: $d$-dependence of the difference between the numerical and analytic values of $W_{2}^{2}$, defined as the numerical value minus the analytic one..
+In order to briefly confirm the accuracy of the derived 2-Wasserstein distance, we compare the analytical expression given in Eq.
+(10) with the original integral expression in Eq.
+(8) for a simple one-dimensional Gaussian case.
+We set the standard deviation to $\sigma=1$.
+The numerical evaluation of Eq.
+(8) is performed using the trapezoidal rule with uniform discretization.
+The integration range is restricted to $q\in[-5\sigma,5\sigma]$, and the representative point $q^{\prime}_{k}$ of each discretization cell is chosen as its midpoint.
+The numerical results shown in Fig.
+1 indicate that the derived analytic expression for $W_{2}$ accurately reproduces the original integral form, particularly in the small-$d$ regime.